# The Art of Saying "Maybe": A Conformal Lens for Uncertainty Benchmarking in VLMs

### v2

@@ Abstract
Vision-Language Models (VLMs) have achieved remarkable progress in complex visual understanding across scientific and reasoning tasks. While performance benchmarking has advanced our understanding of these capabilities, the critical dimension of uncertainty quantification has received insufficient attention. Therefore, unlike prior conformal prediction studies that focused on limited settings, we conduct a comprehensive uncertainty benchmarking study, evaluating 16 state-of-the-art VLMs (open and closed-source) across 6 multimodal datasets with 3 distinct scoring functions. Our findings demonstrate that larger models consistently exhibit better uncertainty quantification; models that know more also know better what they don’t know. More certain models achieve higher accuracy, while mathematical and reasoning tasks elicit poorer uncertainty performance across all models compared to other domains. This work establishes a foundation for reliable uncertainty evaluation in multimodal systems.
The Art of Saying "Maybe": A Conformal Lens for Uncertainty Benchmarking in VLMs
Asif Azad1, Mohammad Sadat Hossain1, MD Sadik Hossain Shanto1, M Saifur Rahman1, Md Rizwan Parvez2 1Bangladesh University of Engineering & Technology (BUET) 2Qatar Computing Research Institute (QCRI) Correspondence: asifazad0178@gmail.com
@@ 2 Related Works
Uncertainty quantification has long been a focus of machine learning (Abdar et al., 2021), particularly for applications involving risk-sensitive decision-making. Classical techniques span Bayesian neural networks (Blundell et al., 2015), deep ensembles (Lakshminarayanan et al., 2017), and calibration-based methods (Guo et al., 2017). While these approaches have proven effective in low-dimensional settings, they are often computationally prohibitive or insufficiently expressive for deep multimodal models.
Conformal prediction is a well-established uncertainty quantification method that offers statistical guarantees and has been successfully applied across various domains (Zhou et al., 2025). Its distribution-free, model-agnostic, and computationally efficient nature makes it particularly suitable for large-scale models. Recent work has applied conformal prediction to LLMs (Angelopoulos and Bates, 2021; Ye et al., 2024) and VLMs (Kostumov et al., 2024), providing coverage guarantees via prediction sets. However, prior studies were limited to text-only models or evaluated VLMs on simpler benchmarks with outdated model selections. Our work extends this paradigm by incorporating complex reasoning tasks and systematically evaluating a comprehensive, up-to-date collection of state-of-the-art VLMs across diverse multimodal contexts.
The emergence of VLMs has shifted attention toward multimodal understanding. Evaluation benchmarks have evolved accordingly, focusing on various aspects of performance including visual reasoning (Zellers et al., 2019), hallucination detection (Liu et al., 2022; Sadat et al., 2023; Islam et al., 2024), and multimodal knowledge (Xu et al., 2023).
Beyond performance evaluation, several parallel research directions address related challenges such as hallucination detection (Rawte et al., 2023), interpretability (Bommasani et al., 2022), and model confidence estimation (Hendrycks and Gimpel, 2018). These approaches provide valuable insights but often lack the formal statistical guarantees required for rigorous uncertainty quantification. In contrast, our work leverages conformal prediction, which offers a unified, distribution-free framework that is both computationally feasible for large-scale foundation models and accompanied by theoretical coverage guarantees. This makes it particularly advantageous compared to heuristic-based methods for hallucination detection or post-hoc interpretability techniques, positioning conformal prediction as a principled path forward for benchmarking uncertainty in VLMs.
@@ 3 Conformal Prediction
Conformal prediction provides a statistically rigorous, distribution-free framework for uncertainty quantification. It constructs prediction sets that contain the true output with a specified probability. For any model $f$ that maps an input $X$ to a probability distribution over a finite label space $Y$, conformal prediction constructs a prediction set $C(X)\subseteq Y$ such that:
@@
where $\alpha$ is the desired error rate.
To construct these sets, one defines a score function $s(X,y)$, which reflects the incompatibility between input $X$ and label $y$. The prediction set is then constructed through the following procedure:
@@
    Compute conformal scores $s_{i}=s(X_{i}^{\text{cal}},Y_{i}^{\text{cal}})$ for each example in a held-out calibration set $D_{\text{cal}}=\{(X_{1}^{\text{cal}},Y_{1}^{\text{cal}}),\ldots,(X_{n}^{\text{cal}},Y_{n}^{\text{cal}})\}$.
@@
    Calculate a threshold $\hat{q}$ as the $\lceil(n+1)(1-\alpha)\rceil/n$ quantile of these calibration scores:
@@
    For any test input $X$, construct the prediction set by including all labels with scores not exceeding the threshold:
@@
In practice, conformal prediction relies on heuristic scoring functions, and different choices can lead to different behaviors in calibration and set size. In our study, we focus on three such scoring functions - Least Ambiguous Classifier (LAC), Adaptive Prediction Sets (APS), and Marginal Score (MS), and systematically compare their behaviors across tasks and models. To further validate our findings, we also analyze entropy-based uncertainty as a complementary measure.
Three principal scoring functions are commonly used in conformal prediction for classification:
@@
where $f(X)_{y}$ denotes the model’s predicted probability for class $y$. This approach penalizes low-confidence predictions, assigning higher scores to less likely labels and thus favoring more confident predictions in the conformal set construction.
@@
which sums the probabilities of all classes with at least as much support as $y$, effectively incorporating the model’s ranking of classes. APS adapts the conformal set size to the ambiguity present in the predictions, making it particularly useful when the model’s probability distribution is diffuse.
@@
where $f(X)_{(1)}$ and $f(X)_{(2)}$ are the top-1 and top-2 predicted probabilities, respectively. This score captures the model’s decisiveness by directly measuring the confidence gap between the most likely and second most likely classes, making it a natural fit for high-ambiguity tasks where subtle distinctions matter.
@@ 4 Datasets
We evaluate uncertainty in VLMs across six diverse, challenging datasets, each probing different aspects of multimodal reasoning and understanding:
@@
The Massive Multi-discipline Multimodal Understanding (MMMU) dataset (Yue et al., 2024a) is a large-scale benchmark designed to assess VLMs on college-level, expert-written questions spanning 30 disciplines, including science, medicine, engineering, and the humanities.
@@
MMMU-Pro (Yue et al., 2024b) is an extension of MMMU, curated to provide more challenging and professionally oriented questions. It emphasizes real-world scenarios and domain-specific expertise, increasing the complexity of both the visual and textual components.
@@
ScienceQA (Lu et al., 2022) is a multimodal benchmark focused on elementary and middle school science questions. It contains over 21,000 questions covering natural sciences, physics, and biology, many of which are accompanied by images such as diagrams or illustrations. The dataset tests the model’s ability to integrate visual information with scientific knowledge.
@@
The AI2 Diagrams (AI2D) dataset (Kembhavi et al., 2016) consists of over 15,000 elementary science diagram questions. Each question is paired with a labeled diagram and multiple-choice answers, requiring the model to interpret visual elements, spatial relationships, and scientific concepts depicted in the diagrams.
@@
MathVision (Wang et al., 2024, 2025) is a visual math reasoning benchmark that presents mathematical problems embedded in images, such as graphs, geometric figures, or handwritten equations. The dataset evaluates the model’s ability to extract quantitative information from visuals and perform mathematical reasoning.
@@
WorldMedQAV (Matos et al., 2024) is a medical visual question answering dataset featuring clinical images (e.g., X-rays, pathology slides) and expert-authored multiple-choice questions. It is designed to assess VLMs’ capabilities in medical image interpretation and diagnostic reasoning, reflecting real-world healthcare scenarios.
Together, these datasets provide a comprehensive testbed for evaluating uncertainty in VLMs across a spectrum of domains, modalities, and reasoning challenges.
@@ 5.1 Prompting
Our prompting strategy employed a three-part structure across all datasets. First, we used dataset-specific system messages that established the VLM’s role (e.g., "scientific diagram analyzer" for AI2D, "medical image diagnostician" for WorldMedQAV). These messages oriented the model to the domain context while maintaining consistent instruction patterns.
Second, we included zero-shot task instructions that briefly described the upcoming question type without revealing solving strategies. For instance, MathVision prompts began with "I will show you an image along with a multiple-choice math question." This framing provided context without biasing model responses.
Finally, all prompts concluded with a standardized instruction directing models to "Only respond with the option letter" to ensure consistent output format for uncertainty analysis. This standardized approach eliminated prompt variability as a confounding factor in our experiments. Complete prompts are provided in Appendix A.
5.2 Inference Setup
We implemented a carefully controlled inference pipeline across different computing platforms based on model size and availability. For all small models ($\leq$ 7B parameters), we conducted inference on P100 and T4 GPUs via the Kaggle platform.
For our selected VLMs, we prioritized using OpenRouter’s API service whenever available, regardless of model size. This approach covered both large and mid-sized models with API endpoints. For mid-sized models without OpenRouter API availability, we utilized A1000 GPUs on the Runpod platform for efficient inference.
All models processed on Kaggle and Runpod were loaded directly from their official Hugging Face repositories to ensure we used canonical model versions. Throughout all inference methods, we set do_sample=False to employ greedy decoding, making temperature, top-k, and top-p parameters irrelevant to our experimental design.
For each model response, we extracted the log-probabilities assigned to the answer option letters (e.g., A, B, C, D) by examining the token-level scores corresponding to the model’s final output. Since all tasks were multiple-choice and responses were constrained to a single letter, we retrieved the log-probability of the token representing the predicted answer. These probability distributions formed the foundation for our uncertainty quantification through conformal prediction. We implemented conformal prediction with miscoverage rate $\alpha=0.1$ (ensuring 90% coverage probability) and allocated 50% of each dataset for calibration and 50% for testing.
5.3 Evaluated Models
We evaluate 16 vision-language models representing diverse architectures and scaling properties. Our open-source selection includes Llama-4-Scout, Gemma-3 (Team et al., 2025) (4B/12B/27B), InternVL3 (Zhu et al., 2025) (1B/2B/8B), Molmo variants (Deitke et al., 2024) (1B/7B), Qwen2.5-VL (Bai et al., 2025) (3B/72B), Llava-1.5 (Liu et al., 2024) (7B/13B), and Pixtral (Agrawal et al., 2024)(12B). For proprietary models, we test GPT-4.1-nano and GPT-4o-mini, the only commercial VLMs providing token probabilities required for conformal analysis.
This meticulate choice of models allows controlled comparisons of uncertainty characteristics across model sizes, architectures, and development paradigms (open/closed). We exclude other proprietary models (e.g., Gemini, Claude ) due to their API limitations on probability access, which is essential for our conformal prediction framework.
5.4 Evaluation Metrics
Our primary uncertainty quantification (UQ) metric is Set Size (SS), which measures the average size of conformal prediction sets:
$\text{SS}=\frac{1}{|D_{test}|}\sum_{(X_{t},Y_{t})\in D_{test}}|C(X_{t})|$ (7)
A smaller SS indicates more precise uncertainty estimates, with SS=1 representing perfect certainty when the prediction is correct. We complement this with traditional Accuracy (Acc) to assess prediction correctness:
Finally, we verify the statistical guarantee of our conformal framework through Coverage Rate (CR):
CR must maintain at least $(1-\alpha)$ coverage across all test cases. Together, these metrics (SS, Acc, CR) evaluated across LAC, MS, and APS score functions provide a complete assessment of both prediction quality and uncertainty reliability.
6 Results
@@ 6.1 Uncertainty Performance Analysis
Table 2 shows set sizes across our evaluated VLMs and conformal scoring functions. LAC scoring produces the smallest set sizes across most models, indicating its effectiveness for uncertainty quantification in vision-language tasks. The consistent pattern of smaller set sizes for larger models within each family (e.g., Qwen-VL 72B vs. 3B) demonstrates that scaling benefits uncertainty calibration.
Figure 1 demonstrates the strong inverse relationship between accuracy and set size, confirming that more accurate models generally produce more concentrated prediction sets with better calibration. As shown in Figure 8, this negative correlation represents a fundamental principle: models that perform better also express more appropriate confidence levels.
Figure 2 shows the distribution of set sizes across different uncertainty methods. We can see that APS scheme demonstrates the least variability.
Figure 2: Set size distribution across different uncertainty methods.
@@ 6.2 Accuracy Performance Analysis
Table 1 reveals clear performance patterns: (1) larger models within the same family consistently achieve higher accuracy; and (2) significant performance gaps exist between datasets, with MMMU-Pro being most challenging (20.85% average accuracy) and ScienceQA most approachable (73.78%). Among open-source models, InternVL 8B delivers the strongest performance, particularly excelling on AI2D and ScienceQA tasks.
@@
Figure 5: Coverage rates across datasets and models. All scoring methods maintain at least 90% coverage in most cases.
Figure 5 shows that model size correlates positively with accuracy and inversely with set size. Models cluster by parameter count, with larger models (>10B) consistently occupying the upper-left region of higher accuracy and smaller set sizes, demonstrating that scaling improves both performance and uncertainty calibration.
6.3 Coverage Rate Analysis
Table 3 and Figure 5 verify that our conformal prediction framework achieves at least $(1-\alpha)=90\%$ coverage in most cases, validating its reliability. The few instances where coverage falls slightly below target show minimal deviation. Overall, coverage is most challenging to maintain for complex reasoning tasks like MathVision and MMMU-Pro, though the framework still performs robustly across all domains.
6.4 Model-Specific Uncertainty Performance
@@
Figure 6 reveals distinct "uncertainty signatures" for each model family. From the figure, it is evident across all model families (Gemma, Qwen-VL, InternVL) that bigger variants demonstrate better confidence in terms of uncertainty quantification while maintaining the specified coverage rate.
Figure 7 provides a comprehensive comparison of uncertainty performance metrics across all evaluated model families. InternVL demonstrates superior uncertainty quantification capabilities, surpassing all competitors in this critical dimension. Llama-4-Scout follows as the second-best performer, suggesting potential architectural advantages in these two model families that may contribute to their enhanced calibration and confidence estimation.

### v3

# The Art of Saying “Maybe”: A Conformal Lens for Uncertainty Benchmarking in VLMs

@@ Abstract
Vision-Language Models (VLMs) have achieved remarkable progress in complex visual understanding across scientific and reasoning tasks. While performance benchmarking has advanced our understanding of these capabilities, the critical dimension of uncertainty quantification has received insufficient attention. Therefore, unlike prior conformal prediction studies that focused on limited settings, we conduct a comprehensive uncertainty benchmarking study, evaluating 18 state-of-the-art VLMs (open and closed-source) across 6 multimodal datasets with 3 distinct scoring functions. For closed-source models lacking token-level logprob access, we develop and validate instruction-guided likelihood proxies. Our findings demonstrate that larger models consistently exhibit better uncertainty quantification; models that know more also know better what they don’t know. More certain models achieve higher accuracy, while mathematical and reasoning tasks elicit poorer uncertainty performance across all models compared to other domains. This work establishes a foundation for reliable uncertainty evaluation in multimodal systems.
The Art of Saying “Maybe”: A Conformal Lens for Uncertainty Benchmarking in VLMs
[…]
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 2 Related Works
Uncertainty quantification has long been central to machine learning, especially for risk-sensitive decisions (Abdar et al., 2021). Classic methods include Bayesian neural networks (Blundell et al., 2015), deep ensembles (Lakshminarayanan et al., 2017), and calibration techniques (Guo et al., 2017). These work well in low dimensions but can be computationally heavy or lack expressiveness for deep multimodal models.
Conformal prediction provides robust, statistical guarantees and applies across domains (Zhou et al., 2025). It’s distribution-free, model-agnostic, and efficient, making it ideal for large-scale models. Recent efforts have used it for LLMs (Angelopoulos and Bates, 2021; Ye et al., 2024) and VLMs (Kostumov et al., 2024), delivering coverage via prediction sets. Still, past work focused on text-only models or simpler benchmarks with outdated VLMs. In contrast, while Kostumov et al. (2024) mainly examined accuracy–uncertainty alignment and scaling effects, our study investigates a broader set of five research questions covering scoring efficiency, calibration–accuracy correlations, architecture and scale influences, black-box model applicability, and domain-specific trends. We also evaluate a larger and more diverse suite of state-of-the-art proprietary and open VLMs across multimodal reasoning tasks, extending beyond alignment analysis toward a unified, theoretically grounded uncertainty benchmarking framework.
VLMs have spotlighted multimodal understanding, with benchmarks targeting visual reasoning (Zellers et al., 2019), hallucination detection (Liu et al., 2022; Sadat et al., 2023; Islam et al., 2024), and multimodal knowledge (Xu et al., 2023).
Related research covers hallucination (Rawte et al., 2023), interpretability (Bommasani et al., 2022), and confidence estimation (Hendrycks and Gimpel, 2018). These offer insights but often miss formal guarantees for uncertainty quantification. Our approach uses conformal prediction for a unified, distribution-free framework that’s practical for big foundation models and backed by theory. It outperforms heuristic hallucination detectors or post-hoc interpretability methods, paving a solid way to benchmark VLM uncertainty.
@@ 3 Conformal Prediction
Conformal prediction offers a statistically rigorous, distribution-free framework for uncertainty quantification. It builds prediction sets that include the true output with a specified probability. For any model $f$ mapping input $X$ to a probability distribution over finite label space $Y$, it creates $C(X)\subseteq Y$ such that:
@@
where $\alpha$ is the error rate.
To build these sets, use a score function $s(X,y)$ that measures incompatibility between $X$ and $y$. The process is:
@@
    Compute scores $s_{i}=s(X_{i}^{\text{cal}},Y_{i}^{\text{cal}})$ for calibration set $D_{\text{cal}}=\{(X_{1}^{\text{cal}},Y_{1}^{\text{cal}}),\ldots,(X_{n}^{\text{cal}},Y_{n}^{\text{cal}})\}$.
@@
    Find threshold $\hat{q}$ as the $\lceil(n+1)(1-\alpha)\rceil/n$ quantile of scores:
@@
    For test input $X$, form $C(X)=\{y\in\mathcal{Y}:s(X,y)\leq\hat{q}\}$:
@@
In practice, it uses heuristic scoring functions, which influence calibration and set size. We examine three: Least Ambiguous Classifier (LAC), Adaptive Prediction Sets (APS), and Marginal Score (MS), comparing them across tasks and models. We also check entropy-based uncertainty as a complement.
Common scoring functions for classification include:
@@
where $f(X)_{y}$ is the predicted probability for $y$. It penalizes low-confidence predictions, favoring confident ones in sets.
@@
summing probabilities of classes as or more supported than $y$. It adjusts set size to prediction ambiguity, suiting diffuse distributions.
@@
where $f(X)_{(1)}$ and $f(X)_{(2)}$ are top-1 and top-2 probabilities. It measures confidence gap, ideal for high-ambiguity tasks.
@@ 4 Datasets
We test VLM uncertainty on six diverse, challenging datasets probing multimodal reasoning:
@@
The Massive Multi-discipline Multimodal Understanding (MMMU) dataset (Yue et al., 2024a) assesses VLMs on college-level questions across 30 disciplines like science, medicine, engineering, and humanities.
@@
MMMU-Pro (Yue et al., 2024b) extends MMMU with tougher, professional questions emphasizing real-world scenarios and domain expertise, ramping up visual and textual complexity.
@@
ScienceQA (Lu et al., 2022) focuses on elementary/middle school science with over 21,000 questions in natural sciences, physics, and biology, often with images like diagrams. It tests visual-scientific integration.
@@
The AI2 Diagrams (AI2D) dataset (Kembhavi et al., 2016) has over 15,000 elementary science questions paired with labeled diagrams and multiple-choice answers, demanding interpretation of visuals, spatial relations, and concepts.
@@
MathVision (Wang et al., 2024, 2025) is a visual math benchmark with problems in images like graphs or equations, evaluating extraction of quantitative info and math reasoning.
@@
WorldMedQAV (Matos et al., 2024) features clinical images (e.g., X-rays, slides) with expert multiple-choice questions, testing medical image interpretation and diagnostic reasoning in healthcare contexts.
These datasets create a broad testbed for VLM uncertainty across domains, modalities, and reasoning types.
@@ 5.1 Prompting
We used a three-part prompting strategy across all datasets. First, dataset-specific system messages set the VLM’s role, like “scientific diagram analyzer” for AI2D or “medical image diagnostician” for WorldMedQAV. These guided models to the domain while keeping instructions consistent.
Second, zero-shot task instructions gave a quick overview of the question type, without hints on solving. For example, MathVision prompts started with “I will show you an image along with a multiple-choice math question.” This setup offered context but avoided biasing responses.
Finally, every prompt ended with “Only respond with the option letter” to standardize outputs for uncertainty analysis. This kept prompt differences from skewing results. Full prompts are in Appendix A.
@@
5.2 Inference Setup
We ran a controlled inference pipeline, adapting to model size and availability. Small models ($\leq$ 7B parameters) used P100 and T4 GPUs on Kaggle.
For VLMs, we preferred OpenRouter’s API when available, covering large and mid-sized models. For mid-sized ones without API, we turned to A1000 GPUs on Runpod for smooth runs.
Models on Kaggle and Runpod came straight from Hugging Face repositories to use official versions. Across all, we set do_sample=False for greedy decoding, skipping temperature, top-k, and top-p.
From each response, we pulled log-probabilities for answer letters (A, B, C, D) via token-level scores on the final output. With multiple-choice tasks limited to one letter, we grabbed the predicted answer’s log-prob. These distributions drove our conformal prediction at $\alpha=0.1$ (90% coverage), splitting datasets 50/50 for calibration and testing.
5.3 Evaluated Models
We tested 18 vision-language models spanning architectures and scales. Open-source picks include Llama-4-Scout, Gemma-3 (Team et al., 2025) (4B/12B/27B), InternVL3 (Zhu et al., 2025) (1B/2B/8B), Molmo variants (Deitke et al., 2024) (1B/7B), Qwen2.5-VL (Bai et al., 2025) (3B/72B), Llava-1.5 (Liu et al., 2024) (7B/13B), and Pixtral (Agrawal et al., 2024) (12B). Proprietary ones include GPT-4.1-nano and GPT-4o-mini, which were the only commercial VLMs sharing token probabilities for conformal work. Later we extended our work with proxy methods, to include Gemini-2.5-Flash and Claude Haiku-4.5.
This lineup lets us compare uncertainty traits across sizes, builds, and open/closed paradigms. Since models like Gemini and Claude do not expose token-level probabilities through their APIs, we developed an instruction-guided likelihood proxy (Section 6.7) to extend our evaluation to these closed-source VLMs.
5.4 Evaluation Metrics
Our main UQ metric is Set Size (SS), averaging conformal set sizes:
$\text{SS}=\frac{1}{|D_{test}|}\sum_{(X_{t},Y_{t})\in D_{test}}|C(X_{t})|$ (7)
Lower SS means sharper uncertainty, with SS=1 for perfect certainty on correct predictions. We pair it with Accuracy (Acc) for prediction performance:
We also check Coverage Rate (CR) to confirm guarantees:
CR should hit at least $(1-\alpha)$ overall. These metrics (SS, Acc, CR) across LAC, MS, and APS give a full picture of prediction quality and uncertainty trustworthiness.
6 Results
@@ 6.1 Uncertainty Performance Analysis
Table 2 presents set sizes for our VLMs and conformal scoring functions. LAC yields the smallest sets for most models, highlighting its strength in vision-language uncertainty quantification. Larger models in each family (e.g., Qwen-VL 72B vs. 3B) show consistently smaller sets, underscoring scaling’s role in better calibration.
Figure 1 illustrates the strong inverse link between accuracy and set size: top performers create tighter, better-calibrated sets. As detailed in Appendix C, Figure 9, this negative correlation is a key principle, stronger models reflect appropriate confidence.
Set size distributions across methods appear in Appendix Figure 6, where APS shows the lowest variability.
@@ 6.2 Accuracy Performance Analysis
Table 1 highlights key trends: (1) larger family members hit higher accuracy; (2) datasets vary widely, with MMMU-Pro the toughest (20.85% average) and ScienceQA the easiest (73.78%). Among open-source models, InternVL 8B leads, shining on AI2D and ScienceQA.
Figure 4 confirms model size boosts accuracy and shrinks set sizes. Larger models (>10B) cluster in the high-accuracy, low-set-size zone, proving scaling enhances performance and calibration.
6.3 Coverage Rate Analysis
Coverage validation is in Appendix Table 13 and Figure 8. Our framework meets $(1-\alpha)=90\%$ coverage in most cases, with minor slips. It’s toughest on complex tasks like MathVision and MMMU-Pro, yet holds up across domains.
6.4 Model-Specific Uncertainty Performance
Figure 5 uncovers unique “uncertainty signatures” per family. Across Gemma, Qwen-VL, and InternVL, bigger variants show superior confidence calibration at the target coverage.
Figure 3 compares all families: InternVL tops uncertainty quantification, outpacing rivals. Llama-4-Scout ranks second, hinting at architectural edges for calibration in these lines.